OpenStudy (anonymous):

integrate (lnx)(x^-(theta +1))dx from 1 to infinity

OpenStudy (anonymous):

check your question

OpenStudy (anonymous):

what u mean?

OpenStudy (anonymous):

\[is it \int\limits x ^{-\theta+1}dx\]

OpenStudy (anonymous):

\[\theta \int\limits_{1}^{\infty} (lnx)(x)\]

OpenStudy (anonymous):

with x to the power of -(theta+1) dx

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

\[\int\limits_{1}^{\infty}\ln x \left( x \right)^{-\theta-1}dx\] \[=[\ln x \frac{ x ^{-\theta} }{- \theta }] 1\to \infty -\int\limits_{1}^{\infty} \frac{ 1 }{ x }\frac{ x ^{-\theta} }{ -\theta }dx\] \[=0+\frac{ 1 }{ \theta }\int\limits_{1}^{\infty}x ^{-\theta-1}dx=\frac{ 1 }{ \theta }\left[ \frac{ x ^{-\theta} }{-\theta } \right] 1 \to\infty \] \[=\frac{ -1 }{ \theta ^{2} }\left( \frac{ 1 }{\infty }-\frac{ 1 }{1 } \right)=\frac{ -1 }{\theta ^{2} }\left( 0-1 \right)\] \[=\frac{ 1 }{ \theta ^{2} }\]

OpenStudy (anonymous):

what did u use to integrate it?

OpenStudy (anonymous):

integrate by parts. \[\ln 1=0\]

OpenStudy (anonymous):

what are the parts?

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

do u mind just writing out the parts u separated them into?

OpenStudy (anonymous):

\[\ln x is first part ,and x ^{-\left( \theta+1 \right)} second part.\]

OpenStudy (anonymous):

ok great thank you

OpenStudy (anonymous):

\[\int\limits u v dx=u \int\limits v dx-\int\limits u \prime( \int\limits v dx )dx\]

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

yw